OpenStudy (anonymous):

solve the following system of equations. 3x - 2y = 6 6x - 4y = 14

OpenStudy (anonymous):

use the elimination method...can you do that ?

OpenStudy (shubhamsrg):

or multiply first eqn by 2 ,see if that helps ?

OpenStudy (anonymous):

this question is false?

OpenStudy (anonymous):

what do you mean false ?

OpenStudy (anonymous):

2*(3x - 2y) = 12 but second equation says that 6x-4y = 14

OpenStudy (anonymous):

3x - 2y = 6 -->(-2)3x - 2y = 6 6x - 4y = 14 ------------- -6x + 4y = - 12 (result of multiplying by -2) 6x - 4y = 14 -------------add 0 + 0 = 2 0 = 2 (incorrect) The solution to this equation is NO SOLUTION

OpenStudy (anonymous):

(0,0) (6,14) infinitely many solutions no solutions

OpenStudy (anonymous):

@kelliegirl33 i think your right thanks

OpenStudy (anonymous):

if the equation would have equaled 0 = 0, then it would have been infinite solutions, but since the equation equals 0 = 14, then there is no solution :)

OpenStudy (anonymous):

Trevors dad decided to get him a cell phone for his birthday. cell plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. cell best has a plan that costs $30.00 per month plus an additional $0.05 per minute. how many minutes can trevor talk and both cell plans cost the same amount? @kelliegirl33

OpenStudy (anonymous):

m = months x = minutes 24 + .06x = 30 + .05x (combine like terms) .06x - .05x = 30 - 24 .01x = 6 x = 6/.01 x = 600 check.. 24 + .06(600) = 30 + .05(600) 24 + 36 = 30 + 30 60 = 60 (correct) Trevor can talk 600 minutes

OpenStudy (anonymous):

forget about the m = months....I didn't need that

OpenStudy (anonymous):

He can talk 600 minutes per month

OpenStudy (anonymous):

there were 150 freshman and sophomores at the high school musical Reese and her friends attended. freshman tickets were $2.00 each and sophomore tickets were $3.00 each. the receipts for freshman and sophomore tickets totaled $390. how many sophomores bought tickets?

OpenStudy (anonymous):

f = freshmen s = sophomore f + s = 150 --> f = 150 - s 2f + 3s = 390 now sub 150 - s in for f 2f + 3s = 390 2(150 - s) + 3s = 390 300 - 2s + 3s = 390 s = 390 - 300 s = 90 now sub 90 in for s f + s = 150 f + 90 = 150 f = 150 - 90 f = 60 check... 2f + 3s = 390 2(60) + 3(90) = 390 120 + 270 = 390 390 = 390 (correct) The sophomores that bought tickets were 90.

OpenStudy (anonymous):

an airplane flew 3.5 hours with a 40 mph head wind. the return trip with a tail wind of the same speed took 2.5 hours. find the speed of the plane in still air.

OpenStudy (anonymous):

I am not sure how to do these kinds of problems...sorry

OpenStudy (anonymous):

its okay ill post it :) thank for your help though :) would you mind if i asked u for help some other time?

OpenStudy (anonymous):

you can ask me anytime....I just hope I know how to do the problem you asked

OpenStudy (anonymous):

ok tlk to u later :)

OpenStudy (anonymous):

ok...bye for now :)